OpenStudy (anonymous):

HELP, WILL REWARD!! Divide -2x^3+21x^2-20x-12 by (x-3)

OpenStudy (zehanz):

Have you studied long division and/or synthetic division? You'll need one of these techniques...

OpenStudy (anonymous):

No I havent

OpenStudy (zehanz):

How can you answer the question then? Are you supposed to "invent" these methods yourself?

OpenStudy (anonymous):

Nope

OpenStudy (zehanz):

Synthetic division goes like this: First, write down the solution of x - 3 = 0, so that is 3: 3 | (the vertical line is to separate 3 from the numbers that will follow) Then, write down (next to each other) all the coefficients of the complicated part, put a line under it (leaving som space) and drop the first coefficient below the line: 3 | -2 21 -20 -12 ------------------- -2 Now multiply 3 with that -2 and write the result (-6) in the next column. Add. Put 15 in this column, below the line: 3 | -2 21 -20 -12 -6 ------------------- -2 15 Now do the same with the next coefficient: 3*15=45 in the next column; add: 3 | -2 21 -20 -12 -6 45 ------------------- -2 15 25 Now the same thing once more: 3 | -2 21 -20 -12 -6 45 75 ------------------- -2 15 25 63 The first three numbers below the line are the coefficients of the plynomial that is left after the division by x - 3. Because the last number is not 0, but 63, the division has that as remainder.

OpenStudy (zehanz):

All in all, you can write \(\dfrac{-2x^{3}+21x^{2}-20x-12}{x-3}=-2x^{2}+15x+25+\dfrac{63}{x-3}\). That's it, you're done! It all sounds complicated, but all you have to do is practise a little... To be fair, this is of course just a trick. I think you should study some algebra to see what is happening and why this trick works. In the mean time, use it to "solve" this kind of problems!